OpenStudy (anonymous):

The population of a region is growing exponentially. There were 20 million people in 1985 (t = 0) and 62 million in 1995. Find an expression for the population at any time t, in years. Use the general exponential function and remember to use exact values. P(t) = ? (in millions) What population would you predict for the year 2005? What is the doubling time in years? (Round your answer to one decimal place.)

OpenStudy (anonymous):

Okay, so we know it is an exponential function. It should be in the form:\[\large P(t) = ce^{kt}\] First we solve for \(c\). When is \(e^{kt} = 1\)?

OpenStudy (anonymous):

im so confused I have no idea how to solve

OpenStudy (anonymous):

Since \(e^0 = 1\) then it is when \(kt = 0\). Does that make sense?

OpenStudy (anonymous):

yea I get that much

OpenStudy (anonymous):

\(kt = 0\) when \(k=0\) or \(t=0\). We know that \(k\neq 0\). But we do know when \(t=0\).

OpenStudy (anonymous):

\[\Large P(0) = ce^{k\cdot 0} \Rightarrow 20 = ce^{k\cdot 0}\] Can you solve for \(c\) here?

OpenStudy (anonymous):

Also, do you know what I just did?

OpenStudy (anonymous):

i dont know how to solve c

OpenStudy (anonymous):

\[\Large P(0) = ce^{k\cdot 0} \Rightarrow 20 = ce^{k\cdot 0} \Rightarrow 20 = c\cdot e^0 \Rightarrow 20 = c\] So \(c = 20\). What level of math are you in?

OpenStudy (anonymous):

Basically \(c = P(0)\). So \(c\) will always be our initial amount in these exponential functions.

OpenStudy (anonymous):

Thus we have \[P(t) = 20e^{kt}\]Our next step is to solve for k.

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

can you please just solve it for me because im on a timed problem

OpenStudy (anonymous):

We have to use the fact that \[P(10) = 62\]We were given this when they said in 1995 the population was 1995.

OpenStudy (anonymous):

ok i got that.

OpenStudy (anonymous):

how do we get the general form to find any time

OpenStudy (anonymous):

So \[\Large 62 = 20e^{10k} \Rightarrow \frac{62}{20} = e^{10k}\Rightarrow \ln\left(\frac{62}{20}\right) = 10k \Rightarrow k = \frac{1}{10}\ln\left(\frac{62}{20}\right) \]

OpenStudy (anonymous):

whats the final answer

OpenStudy (anonymous):

It doesn't matter to me if you're timed or not.

OpenStudy (anonymous):

p(t)=20(3.1^t/10)???

OpenStudy (anonymous):

No, k = 1.131/10

OpenStudy (anonymous):

So it's \[\Large P(t) = 20e^{0.1131t}\]

OpenStudy (anonymous):

Doubling time \(t_0\) such that \(P(t_0) = 2P(0)\)

OpenStudy (anonymous):

So solve for \(t_0\) in: \[\Large 40 = 20e^{.113t_0}\]

OpenStudy (anonymous):

thank you. i got it

OpenStudy (anonymous):

how would i use that to find the doubling time.

OpenStudy (anonymous):

One variable, one equation. Don't you know how to isolate a variable?

OpenStudy (anonymous):

of p(t)=20(3.1 ^t/10)

OpenStudy (anonymous):

i dont know the doubling time formula